### v1

@@ Abstract
Trustworthiness is essential for the precise and interpretable application of artificial intelligence (AI) in medical imaging. Traditionally, precision and interpretability have been addressed as separate tasks, namely medical image analysis and explainable AI, each developing its own models independently. In this study, for the first time, we investigate the development of a unified medical vision-language pre-training model that can achieve both accurate analysis and interpretable understanding of medical images across various modalities. To build the model, we construct MedConcept-23M, a large-scale dataset comprising 23 million medical image-text pairs extracted from 6.2 million scientific articles, enriched with concepts from the Unified Medical Language System (UMLS). Based on MedConcept-23M, we introduce ConceptCLIP, a medical AI model utilizing concept-enhanced contrastive language-image pre-training. The pre-training of ConceptCLIP involves two primary components: image-text alignment learning (IT-Align) and patch-concept alignment learning (PC-Align). Specifically, IT-Align facilitates the global alignment of medical image and text representations, while PC-Align uses UMLS knowledge for detailed alignment between image patches and their corresponding conceptual representations. This dual alignment strategy enhances the model’s capability to associate specific image regions with relevant concepts, thereby improving both the precision of analysis and the interpretability of the AI system. We conducted extensive experiments on 5 diverse types of medical image analysis tasks, spanning 51 subtasks across 10 image modalities, with the broadest range of downstream tasks. The results demonstrate the effectiveness of the proposed vision-language pre-training model. Further explainability analysis across 6 modalities reveals that ConceptCLIP achieves superior performance, underscoring its robust ability to advance explainable AI in medical imaging. These findings highlight ConceptCLIP’s capability in promoting trustworthy AI in the field of medicine.
@@
Artificial Intelligence, Trustworthiness, Medical Image Analysis, Explainable AI, Image-Language Pre-training, Concept Enhancement
@@ 1 Introduction
Trustworthiness in medical image analysis refers to the accuracy and interpretability of AI systems, which are essential for the effective deployment in clinical settings [1]. Ensuring that AI systems deliver precise and interpretable results is crucial for reliable clinical decision-making. Traditionally, the development of trustworthy medical AI systems involves two distinct components: a medical image analysis module [2] and an explainable AI module [3]. The medical image analysis module is designed to produce outputs that meet clinical requirements, such as providing accurate classification labels for diagnosis or generating coherent textual outputs for report generation and visual question answering. Meanwhile, the explainable AI module audits and elucidates the decision-making process of medical image analysis, offering insights into the model’s decision. Despite advancements in these areas, there remains a gap in achieving both tasks using a unified system.
In this study, we propose that a single medical vision-language pre-training model can effectively address both analysis and explanation tasks across various modalities. This proposition is substantiated by two key aspects. First, vision-language models pre-trained with large-scale image-text pairs have demonstrated promising performance across a wide spectrum of image analysis tasks in both general [4, 5, 6] and medical domains [7, 8]. Second, these models exhibit robust zero-shot classification capabilities [4], which provide a robust foundation for zero-shot concept annotation. As stated in  [3], stronger zero-shot concept annotation ability could result in higher-quality explanations within AI systems. Thus, this capability serves as a bridge, motivating the integration of analysis and explanation modules in a unified framework.
Figure 1: Overview of ConceptCLIP and MedConcept-23M. (a) Bar plot depicting the number of image-text pairs for each image type in MedConcept-23M. (b) Evaluation tasks categorized by modality and subtask, along with the number of images in each subtask. (c) ConceptCLIP demonstrates state-of-the-art performance across various medical imaging tasks. (d) Pre-training strategy for ConceptCLIP, involving joint pre-training with global image-caption alignment and local patch-concept alignment. (e) Examples of tasks that ConceptCLIP is capable of handling.
Implementing a vision-language pre-training model capable of handling various modalities offers several key advantages for enhancing trustworthiness in medical AI systems. First, such a model simplifies the construction of these systems by integrating two traditionally separate modules—analysis and explainability—into a single framework. Second, it addresses the challenge of explanation data [9] scarcity, which is a significant barrier in developing explainable AI systems. Traditionally, creating explanations for specific types of medical images requires substantial resources. This includes getting help from medical experts in different fields to annotate explanations for these images [10, 9, 11, 12, 13]. In scenarios where there are few explanation annotations available, conducting explainable AI becomes nearly impossible. However, a robust vision-language model capable of operating across diverse modalities can significantly alleviate this issue by enabling high-quality zero-shot concept annotation. This capability makes explainable AI feasible even in medical domains with limited explanation annotations, thereby advancing the development of trustworthy medical AI systems.
@@
To support this endeavor, we introduce MedConcept-23M, a large-scale dataset comprising 23 million image-text-concept triples derived from 6.2 million scientific articles and enriched with medical concepts from the Unified Medical Language System (UMLS) [14]. Unlike previous datasets [7, 8], MedConcept-23M is enhanced with the knowledge of medical concept, providing fine-grained textual information paired with various image modalities. As shown in Figure 1 (a), MedConcept-23M includes a diverse range of image types within the field of biomedicine, such as radiology, microscopy, photography, 3D reconstructions, and signal images, covering a broad spectrum of visual data.
Building on this dataset, we develop ConceptCLIP, the first medical vision-language pre-training model designed for both accurate analysis and interpretable understanding of medical images across multiple modalities. As shown in Figure 1 (d), given the nuanced information inherent in medical images, which is critical for precise diagnosis and treatment, ConceptCLIP employs two main pre-training components: image-text alignment (IT-Align) and patch-concept alignment (PC-Align). IT-Align facilitates the global alignment of medical image and text representations, while PC-Align refines this alignment by incorporating concepts from UMLS [14], enabling local alignment between image patches and their corresponding concepts. This approach enhances the precision of medical image analysis by combining global and fine-grained local alignment across image and text modalities.
@@
To evaluate the performance of ConceptCLIP, we develop the most comprehensive benchmark on 5 types of medical image analysis tasks, spanning 51 medical image analysis subtasks across 10 image modalities (Figure 1 (b)). Our evaluation includes 36 subtasks for medical image diagnosis, incorporating zero-shot, linear probes, and full fine-tuning settings. Additionally, it includes 2 subtasks for cross-modal retrieval, 2 subtasks for report generation, 2 subtasks for visual question answering, and 9 subtasks for whole-slide image analysis. As depicted in Figure 1, ConceptCLIP consistently achieves state-of-the-art results. The outcomes in cross-modal retrieval, zero-shot classification, and visual question answering demonstrate that ConceptCLIP is the leading vision-language pre-training model in the medical domain. The results from linear probes, full fine-tuning classification, medical report generation, and whole-slide image analysis underscore the robustness of its vision encoder, highlighting its capability and potential applications across various medical imaging modalities. Further explainability analysis across 6 modalities (Figure 1(b) and Figure 1(c)) confirms ConceptCLIP superior performance, highlighting its strong potential to advance explainable AI in medical imaging and enhance the trustworthiness of AI in medicine.
In summary, our contributions are threefold:
@@
  • We curate MedConcept-23M, a large-scale medical dataset comprising 23 million medical image-text pairs derived from 6.2 million PubMed articles. Utilizing the UMLS system, each image-text pair is supplemented with UMLS concept information, resulting in a fine-grained knowledge-enhanced dataset with 23 million image-text-concept triplets.
@@
  • We develop ConceptCLIP, the state-of-the-art medical vision-language pre-training model capable of accurate analysis and interpretable understanding of medical images across diverse modalities, leveraging both image-text and patch-concept alignments.
@@
  • We conduct the most extensive evaluation on 5 types of tasks spanning 51 medical image analysis subtasks, across 10 image modalities. The results demonstrate that ConceptCLIP achieves state-of-the-art performance, establishing it as the leading vision-language pre-training model in the medical domain and featuring the most advanced vision encoder, which has potential applications across various image modalities. Further explainability analysis across 6 modalities highlights the superior performance of the proposed model, showcasing the strong capability to advance trustworthy AI in the field of medicine.
@@ Results
Large-scale pre-training dataset with image-text-concept triplets and comprehensive evaluation benchmark
This section outlines the creation and application of MedConcept-23M, a pioneering dataset that serves as the foundational pre-training resource for ConceptCLIP, a state-of-the-art medical vision-language model. Additionally, we detail the comprehensive evaluation benchmark employed to assess ConceptCLIP across a diverse array of downstream tasks.
Overview of the MedConcept-23M Dataset
The MedConcept-23M dataset is designed to facilitate the pre-training of ConceptCLIP, offering a large-scale resource of 23 million medical image-text pairs, each augmented with UMLS concept information. Specifically, derived from the extensive PubMed Central Open Access Subset (PMC-OA) [15], MedConcept-23M provides a rich treasure of medical knowledge across various domains. By leveraging advanced concept extraction techniques [^1] https://github.com/allenai/scispacy, each image-text pair is enriched with relevant UMLS concepts, resulting in a dataset that is both voluminous and semantically rich. This comprehensive dataset forms the backbone for pre-training ConceptCLIP, enabling it to capture intricate medical relationships and semantics.
@@
Following its pre-training on MedConcept-23M, ConceptCLIP is subjected to a rigorous evaluation across 51 distinct medical image analysis subtasks, as detailed in Table LABEL:tab:datasets. This benchmark is the most exhaustive to date, encompassing 10 different image modalities and a wide range of tasks, including medical image diagnosis, cross-modal retrieval, report generation, visual question answering, and whole-slide image analysis. Importantly, there is no overlap between MedConcept-23M and the datasets used for evaluation, ensuring an unbiased assessment of ConceptCLIP’s capabilities.
The evaluation benchmark measures the performance of ConceptCLIP across various medical image analysis tasks. Through explainable analyses conducted across six modalities, we highlight ConceptCLIP’s ability to provide interpretable insights, reinforcing its utility in developing trustworthy AI systems in medicine.
Superior diagnostic capabilities of ConceptCLIP in medical image analysis
@@
In this section, we present a comprehensive analysis of the performance of ConceptCLIP in medical image diagnosis across various imaging modalities and experimental settings. The results demonstrate the superior diagnostic capabilities of ConceptCLIP, highlighting its potential as a robust tool in medical image classification.
Zero-shot classification performance
In this experiment, we conduct a systematic evaluation of ConceptCLIP’s ability in the zero-shot setting, where our model is tested on classifying novel classes without requiring additional fine-tuning. Table 3 illustrates the AUC scores for zero-shot classification across different medical imaging modalities, including X-Ray, Fundus, Pathology, Endoscopy, Mammography, Dermoscopy, Ultrasound, OCT, MRI, and CT. ConceptCLIP consistently achieves the highest AUC scores across most datasets, outperforming competing models such as CLIP, SigLIP-400M, PMC-CLIP, and BiomedCLIP. Notably, in the X-Ray modality, ConceptCLIP achieves an average AUC score of 70.82, significantly surpassing the second-best model, BiomedCLIP, which scores 61.05. This trend of superior performance is consistent across other modalities, underscoring the model’s effectiveness in zero-shot settings.
Linear probes with varying training data
To further explore the capabilities of the image encoder of ConceptCLIP, we take 36 subtasks in medical image analysis to evaluate the ability of image representation. The experimental results are demonstrated in the linear probes setting with 1%, 10%, and 100% training data, as detailed in Tables 4, 5, and 6, respectively. ConceptCLIP consistently outperforms other models across all modalities and training data percentages. For instance, with 1% training data, ConceptCLIP achieves an average AUC of 67.95 in the X-Ray category, compared to the second-best average of 64.75 by SigLIP-400M. This performance advantage is maintained and even amplified as the percentage of training data increases, demonstrating the model’s ability to leverage additional data effectively to improve diagnostic accuracy.
Full fine-tuning performance
In the full fine-tuning setting, as depicted in Table 7, ConceptCLIP continues to exhibit outstanding performance across all imaging modalities. The model achieves near-perfect scores in several datasets, such as the Breast Tumor MRI and Brain Tumor CT, where it reaches an AUC of 100.00. This indicates ConceptCLIP’s capability to fully exploit the available data, achieving exceptional diagnostic precision and reliability.
ConceptCLIP shows strong performances in advanced medical tasks
@@
This section delves into the capabilities of ConceptCLIP beyond traditional medical image diagnosis, exploring its performance in cross-modal retrieval, visual question answering, medical report generation, and whole-slide image analysis. The results underscore ConceptCLIP’s versatility and effectiveness across a range of complex tasks, highlighting its potential to enhance various facets of medical data analysis.
Cross-modal retrieval
@@
ConceptCLIP is capable of identifying and retrieving the most relevant textual information given an image input (image-to-text retrieval) and conversely, retrieving images based on a text input (text-to-image retrieval). To assess this capability, we use the PMC-9K and QUILT-1M datasets, as detailed in Tables 8 and 9. ConceptCLIP achieves the highest Recall scores in both Image-to-Text and Text-to-Image tasks. Specifically, in the PMC-9K dataset, ConceptCLIP achieves a Recall@1 of 82.85 for Image-to-Text retrieval, significantly surpassing the second-best model, BiomedCLIP, which scores 73.41. This superior performance demonstrates ConceptCLIP’s proficiency in understanding and linking visual and textual information, a crucial capability for developing integrated medical information systems.
Visual question answering
@@
The visual question answering performance of ConceptCLIP is analyzed using the SLAKE and VQA-RAD datasets. ConceptCLIP achieves the highest overall accuracy in both datasets, with an overall accuracy of 83.91 on SLAKE and 70.78 on VQA-RAD. These results indicate that ConceptCLIP can effectively comprehend and respond to complex medical queries, demonstrating its potential as a decision-support tool in clinical environments where quick and accurate medical decision-making is essential.
Medical report generation
@@
Table 11 presents the performance of ConceptCLIP on the medical report generation task using the MIMIC-CXR and IU X-Ray datasets. ConceptCLIP consistently outperforms or reaches competitive performance with other models, achieving the highest scores in most metrics, including Micro Precision, Micro Recall, Micro F1, BLEU scores, ROUGE-L, METEOR, and CIDEr. Notably, in MIMIC-CXR, ConceptCLIP achieves a Macro F1 score of 4.43 and a BLEU-4 score of 8.36, indicating its superior ability to generate accurate and coherent medical reports. This performance suggests that ConceptCLIP can effectively understand and summarize complex medical information, providing valuable support in clinical documentation.
Whole-slide image analysis
@@
In the realm of whole-slide image analysis, as shown in Table 12, ConceptCLIP demonstrates exceptional performance across various tasks, including cancer diagnosis, molecular subtyping, and survival prediction. For instance, in the Cancer Diagnosis task on the BRACS-3 dataset, ConceptCLIP achieves an AUC score of 91.65, outperforming all other models, especially pathology vision-language models like QuiltNet [16] and PLIP [17]. Similarly, in the Molecular Subtyping task for BRCA, ConceptCLIP records the highest AUC of 74.36. These results highlight ConceptCLIP’s capability to process and analyze high-dimensional pathology images, making it a promising tool for enhancing diagnostic accuracy in histopathology.

### v2

@@ Abstract
The clinical adoption of artificial intelligence (AI) in medical imaging requires models that are both diagnostically accurate and interpretable to clinicians. While current multimodal biomedical foundation models prioritize performance, their black-box nature hinders explaining the decision-making process in clinically meaningful concepts. Here, we present ConceptCLIP, the first explainable biomedical foundation model that achieves state-of-the-art diagnostic accuracy while delivering human-interpretable explanations across diverse imaging modalities. We curate MedConcept-23M, the largest pre-training dataset comprising 23 million image-text-concept triplets across diverse medical modalities, where clinical concepts are derived from the Unified Medical Language System. Leveraging this dataset, we develop ConceptCLIP through a novel dual-alignment approach that simultaneously learns global image-text representations and fine-grained region-concept associations for precise and interpretable medical image analysis. We curate the most extensive evaluation benchmark for multimodal biomedical foundation models, covering 52 clinical tasks spanning 10 imaging modalities. Extensive experiments demonstrate that ConceptCLIP outperforms existing state-of-the-art multimodal biomedical foundation models. Importantly, ConceptCLIP demonstrates superior diagnostic performance while providing human-understandable explanations validated by clinical experts. As the first precise and interpretable biomedical foundation model, ConceptCLIP represents a critical milestone toward the widespread clinical adoption of AI, thereby advancing trustworthy AI in medicine.
@@
Artificial Intelligence, Medical Image Analysis, Explainable AI, Vision-Language Pre-training
@@ 1 Introduction
The clinical adoption of artificial intelligence (AI) systems for medical imaging needs two essential requirements: diagnostic accuracy comparable to specialists and interpretability that supports clinical decision-making [1]. While biomedical foundation models have achieved remarkable accuracy and broad generalizability across diverse imaging modalities [2, 3, 4], they often function as opaque black boxes, lacking fine-grained, clinically meaningful concepts necessary for explanations in medical practice [5]. On the other hand, explainable AI (XAI) methods  [6, 7, 8, 9] provide human-interpretable explanations such as medical concepts to help doctors understand the decision-making process of an AI model. However, these methods are constrained by heavy reliance on expert annotations for a specific clinical setting or modality [10, 11, 12, 13], restricting their broader applicability. The integration of these capabilities—accurate and interpretable medical image analysis with broad generalizability—remains a critical challenge, hindering widespread clinical adoption.
@@
To address this challenge, we develop ConceptCLIP, the first explainable biomedical foundation model that achieves diagnostic performance superior to state-of-the-art foundation models while providing meaningful explanations across diverse clinical settings. Our model is pre-trained on MedConcept-23M, the largest pre-training dataset comprising 23 million image-text-concept triplets derived from 6.2 million scientific articles and enriched with medical terminologies from the Unified Medical Language System (UMLS) [14]. As shown in Fig. 1 (a), MedConcept-23M includes a diverse range of image types within the field of biomedicine, such as pathology, ultrasound, and X-Ray. Unlike previous datasets with only image-text pairs [3, 2], MedConcept-23M is further enhanced with the knowledge of medical concepts (as shown in Fig. 1 (b) and (c)), providing fine-grained textual information paired with various image modalities.
To effectively capture the relationship between medical images and their corresponding concepts, ConceptCLIP employs a dual alignment approach, consisting of image-text alignment (IT-Align) and region-concept alignment (RC-Align). As shown in Fig. 1 (d) and (e), IT-Align facilitates the global alignment of medical image and text representations, while RC-Align enables local alignment between image regions and their corresponding concepts enhanced by knowledge from UMLS [14]. Under this training approach, our model learns to associate medical images with relevant concepts, exploring nuanced clues in the images and concepts that could indicate specific diseases, thereby enhancing the precision of medical image analysis and explainability.
To evaluate the performance of ConceptCLIP, we develop a comprehensive benchmark covering the widest range of tasks, encompassing 52 medical image analysis tasks across 10 image modalities (Fig.  1 (f)). Our evaluation includes 36 tasks for medical image diagnosis, incorporating zero-shot, linear probing, and fully fine-tuning settings, to thoroughly assess ConceptCLIP on medical image diagnosis, the fundamental task in medical imaging, across various modalities. Additionally, we explore the capabilities of ConceptCLIP in other advanced medical tasks, including 15 tasks such as cross-modal retrieval, report generation, visual question answering, and pathology whole-slide image analysis. Furthermore, we conduct extensive explainability analyses across various modalities to evaluate the model’s ability for automatic concept annotation, discovering correlations between concepts and diseases, and employing inherently interpretable models.
@@
As depicted in Fig. 1 (g), Fig. 2, and Fig. 3, ConceptCLIP consistently achieves state-of-the-art results. The outcomes in cross-modal retrieval, zero-shot classification, and visual question answering demonstrate that ConceptCLIP is the leading vision-language pre-training model in the medical domain. Moreover, ConceptCLIP also acts as a well-pretrained vision model in medicine, demonstrating superior performance in linear probing, fully fine-tuning classification, medical report generation, and pathology whole-slide image analysis. The explainability analyses across various modalities (Fig. 1(g) and Fig. 4) confirm the superior performance of ConceptCLIP. As the first explainable foundation model in medicine, ConceptCLIP represents a significant milestone toward clinically trustworthy AI, with broad medical applications (Fig. 1 (h)). This advancement paves the way for the practical deployment of reliable, interpretable AI systems in healthcare settings.
In summary, our contributions are fourfold:
@@
  • We curate MedConcept-23M, the largest pre-training dataset comprising 23 million medical image-text-concept triplets. This dataset is enriched with standardized medical terminologies from the UMLS, making it a fine-grained and knowledge-enhanced resource for developing trustworthy AI in medicine.
@@
  • We develop ConceptCLIP, the first explainable biomedical foundation model via concept-enhanced pre-training. It is capable of achieving state-of-the-art performance while offering an interpretable understanding of medical images across diverse modalities.
@@
  • We conduct an extensive evaluation benchmark of the broadest range across multi-modal biomedical foundation models, including 52 medical image analysis tasks, across 10 image modalities. The results demonstrate that ConceptCLIP achieves state-of-the-art performance among biomedical foundation models, establishing it as the leading medical vision-language pre-training model with strong capability in medical image analysis.
  • We perform comprehensive explainability analysis, exploring how ConceptCLIP enables encoded concept knowledge to provide concept-based explanations, from inherently interpretable model to concept-disease correlation discovery. The results show that ConceptCLIP exhibits promising interpretability, highlighting the model’s strong capability to enhance trustworthy AI in the field of medicine.
Figure 1: Overview of the study. a, Estimated distribution of images in MedConcept-23M dataset by image types. b, Distribution of concepts in each medical image modality studied in this work. c, The wordcloud of concepts in the MedConcept-23M dataset. Word size is proportional to its occurrence in MedConcept-23M. Common nouns and verbs are ignored. Wordclouds for each modality are shown in Extended Data Fig. 5. d, The pre-training process of ConceptCLIP with MedConcept-23M via an image-text alignment (IT-Align) and region-concept alignment (RC-Align) learning. “CUI” refers to “Concept Unique Identifier” in UMLS. “I-T” denotes “image and text”. e, Graphical illustration of the IT-Align in the pre-training process. Clinically related region-concept pairs are gradually learned to align together. f, Statistical information of the evaluation data across different datasets and modalities. g, Comparative performance of ConceptCLIP and other multi-modal biomedical foundation models on various tasks. h, ConceptCLIP can process diverse modalities and perform versatile tasks in medical image analysis and explainable AI, showcasing its outstanding capability in trustworthy AI.
@@ Results
Curating a large-scale medical image-text-concept dataset
The MedConcept-23M dataset is designed to facilitate the pre-training of ConceptCLIP, offering the largest-scale resource of 23 million medical image-text-concept triplets across diverse modalities (Fig. 1 (a)-(c) and Extended Data Fig. 5), each augmented with UMLS terminology information (as described in the Method section). Specifically, derived from the extensive PubMed Central Open Access Subset (PMC-OA) [15], MedConcept-23M provides valuable medical knowledge across various domains. By leveraging advanced concept extraction techniques, each image-text pair is enriched with relevant UMLS concepts, resulting in a dataset that is both voluminous and semantically rich. This comprehensive dataset forms the backbone for pre-training ConceptCLIP, enabling it to capture intricate medical relationships and semantics via a dual-alignment pre-training approach, with a global image-text alignment and a region-concept alignment for pre-training (Fig. 1 (d) and (e)).
@@
Following the pre-training on MedConcept-23M, ConceptCLIP is subjected to a rigorous evaluation across 52 distinct medical image analysis tasks, as detailed in Fig. 1 (f) and Extended Data Table LABEL:tab:datasets. This is the most extensive benchmark for multi-modal medical foundation models, encompassing 10 different image modalities and a wide range of tasks, including medical image diagnosis, cross-modal retrieval, report generation, visual question answering, and pathology whole-slide image analysis. Importantly, there is no overlap between MedConcept-23M and the datasets used for evaluation, ensuring an unbiased assessment of ConceptCLIP’s capabilities. Our evaluation also measures the performance of ConceptCLIP on explainable analyses conducted across various modalities (Fig. 1 (d) and (h)). We highlight ConceptCLIP’s ability to provide plausible explanations, reinforcing its utility in developing trustworthy AI systems in medicine.
Superior diagnostic capabilities of ConceptCLIP across image modalities
@@
In this section, we present a comprehensive analysis of the performance of ConceptCLIP in medical image diagnosis across various imaging modalities in zero-shot, linear probing, and fully fine-tuning settings (Fig. 2). The results demonstrate the superior diagnostic performance of ConceptCLIP, highlighting its outstanding capability in medical image analysis.
ConceptCLIP can perform superior zero-shot diagnosis without task-specific training. In this experiment, we conduct a systematic evaluation of ConceptCLIP’s ability in the zero-shot setting, where our model is tested on classifying novel classes without requiring additional fine-tuning (Fig. 2 (a), (b) and Extended Data Table 3). We evaluate ConceptCLIP on 36 datasets spanning 10 medical imaging modalities in zero-shot settings. In this experiment, ConceptCLIP is compared with other multimodal foundation models in the general domain and medical domain, which can support diverse modalities. By evaluating via AUC (area under the curve) score, our experiment shows that ConceptCLIP consistently outperforms other foundation models across all 10 modalities. For instance, compared to the previous best-performing model trained on diverse medical data, ConceptCLIP achieves absolute AUC gains of 1.46–10.73% ($P<0.001$) across modalities (with an average gain of 6.81%, $P<0.001$), with notable improvements in CT (+10.73%, $P<0.001$), ultrasound (+10.06%, $P<0.001$), and fundus (+8.98%, $P<0.001$). Ablation studies (Extended Data Table 15) reveal two critical drivers: concept-enhanced pre-training (contributing 3.76% of performance gains by incorporating RC-Align in pre-training, $P<0.001$) and top-$K$ local region information for each concept during inference (1.78%, $P<0.001$). This unified framework, combining diverse modalities with concept-level fine-grained medical knowledge, establishes a foundation for generalizable medical artificial intelligence without modality-specific constraints.
ConceptCLIP delivers robust and generalizable medical image representations in linear probing. To further investigate the capabilities of the ConceptCLIP image encoder, we evaluate its performance across 30 datasets in medical image analysis. The efficacy of image representation is assessed using linear probing with 1%, 10%, and 100% of the training data, as depicted in Fig. 2 (a), (c) and detailed in Extended Data Tables 4, 5, and 6. ConceptCLIP consistently surpasses other multi-modal foundation models, which accommodate diverse input modalities, across most modalities and data proportions. For example, with 1% training data, ConceptCLIP achieves an average AUC of 71.48% (95% CI: 62.15% - 80.44%) in the X-Ray category, outperforming the second-best model, SigLIP-400M, which scores 67.73% (95% CI: 59.20% - 76.15%), by 3.75% ($p<0.001$). This trend persists with 10% and 100% training data, where ConceptCLIP demonstrates superior performance among multi-modal foundation models, highlighting its ability to produce
Figure 2: Medical image diagnosis results. a, The graphical illustration depicts the process of zero-shot diagnosis, diagnosis via linear probing, and diagnosis via fully fine-tuning. For zero-shot diagnosis, $\mathbf{t}_{n,\text{cls}}$ and $\mathbf{t}_{n,\text{loc}}$ denote the global text-level and local concept-level representations of class $n$, respectively, while $\mathbf{i}_{m,\text{cls}}$ and $\mathbf{i}_{m,\text{loc}}$ represent the global and local region-level representations of the $m$-th input image (as described in the Method section). b, Zero-shot performance is evaluated using 36 datasets across 10 modalities. Average AUC scores (%) are presented for each specific image modality in the plots. Significance levels at which ConceptCLIP outperforms the best competing method are indicated using a two-sided paired t-test: ***$P<0.001$; **$P<0.01$; *$P<0.5$. c, Linear probing performance is evaluated using 30 datasets across 10 modalities, with the mean ($\pm$±s.d.) reported. Experiments considered 1%, 10%, and 100% proportions of the corresponding training set. d, Fully fine-tuning performance is evaluated using 36 datasets across 10 modalities, with the mean ($\pm$±s.d.) reported.
high-quality image representations for downstream diagnostic tasks across various modalities.
ConceptCLIP improves diagnosis performance in fully fine-tuning. Fully fine-tuning is a fundamental approach for developing AI models in medical image diagnosis. In this setting, as illustrated in Fig. 2 (a), models are fine-tuned on the training set of the specific downstream task prior to evaluation. Fig. 2 (d) and Extended Data Table 7 present the results, where ConceptCLIP consistently demonstrates outstanding performance across all imaging modalities. The model achieves the best performance in 7 out of 10 modalities. In the 3 remaining modalities (endoscopy, dermatology, and pathology), ConceptCLIP exhibits marginally lower performance than SigLIP-400M, with AUC differences of 0.08%, 0.59%, and 0.12%, respectively. These minor gaps may arise because SigLIP-400M’s generic pretraining captures certain low-level visual features more effectively in these domains. However, ConceptCLIP enables superior generalization in most diagnostic settings, as evidenced by its 3.29% AUC gain over SigLIP-400M in X-Ray diagnosis ($P<0.001$). Overall, these findings suggest that in fully fine-tuning, ConceptCLIP significantly outperforms other multi-modal foundation models, indicating its advantages in this context.
ConceptCLIP excels in other versatile medical tasks
@@
This section explores the capabilities of ConceptCLIP beyond medical image diagnosis, examining its performance in cross-modal retrieval, visual question answering, medical report generation, and pathology whole-slide image analysis. The results highlight ConceptCLIP’s versatility and effectiveness across a range of tasks, underscoring its potential to enhance various aspects of medical data analysis.
Figure 3: Results of other medical image analysis tasks. a, This graphical illustration outlines the processes involved in five distinct medical image analysis tasks: text-to-image retrieval, image-to-text retrieval, visual question answering, pathology whole-slide image analysis, and medical report generation. b, Performance in cross-modal retrieval is assessed on two datasets, with Image-to-Text Recall (%) and Text-to-Image Recall (%) reported for each dataset. c, visual question answering tasks are evaluated using accuracy (%) across two datasets, with the mean ($\pm$±s.d.) reported. d, Performance in medical report generation is presented for two datasets using metrics such as BLEU-1,2,3,4, ROUGE-L, METEOR, Micro Precision, Micro Recall, and Micro F1. e, Pathology whole-slide image analysis results are provided across nine tasks within three categories. AUC scores (%) are reported for cancer diagnosis and mutation prediction, while C-index (%) is reported for survival prediction, with the mean ($\pm$±s.d.) reported.
@@
ConceptCLIP showcases strong ability in cross-modal image-text retrieval. ConceptCLIP demonstrates proficiency in image-to-text and text-to-image retrieval, as shown in Fig. 3 (a). The text encoder of ConceptCLIP generates a representation of the input text, which is used to calculate similarities with representations of candidate images encoded by the image encoder. We evaluate this capability using the PMC-9K (we curated it as described in the Method section) and QUILT-1M [16] datasets, comparing ConceptCLIP with other multi-modal foundation models, as detailed in Fig. 3 (b), Extended Data Table 8 and 9. ConceptCLIP achieves the highest Recall scores in both Image-to-Text and Text-to-Image tasks. Specifically, in the PMC-9K dataset, ConceptCLIP attains a Recall@1 of 82.85% (95% CI: 82.05% - 83.60%) for Image-to-Text retrieval, significantly outperforming BiomedCLIP, which scores 73.41% (95% CI: 72.53% - 74.32%), by 9.44% ($P<0.001$). In QUILT-1M, ConceptCLIP achieves a Recall@200 score of 32.50% (95% CI: 31.79% - 33.45%) for Text-to-Image retrieval, representing a 12.88% increase ($P<0.001$) over BiomedCLIP, which scores 19.62% (95% CI: 18.91% - 20.31%). This superior performance demonstrates ConceptCLIP’s proficiency in understanding and linking visual and textual information, a crucial capability for developing integrated medical information retrieval systems.
@@
ConceptCLIP is proficient in medical visual question answering. The visual question answering (VQA) task is crucial for evaluating a model’s ability to comprehend and respond to complex medical queries. As depicted in Fig. 3 (a), a medical image and a question are encoded by the image and text encoders of a foundation model, respectively. These are then processed by a fusion module to calculate the similarity to candidate answers, ultimately selecting the most appropriate answer (described in the Method section). We analyzed the VQA performance of ConceptCLIP using the SLAKE and VQA-RAD datasets. As shown in Fig. 3 (c) and Extended Data Table 10, ConceptCLIP achieves the highest overall accuracy in both datasets, with 83.86% (95% CI: 81.53% - 85.96%) on SLAKE and 70.70% (95% CI: 66.73% - 74.73%) on VQA-RAD, outperforming other multi-modal foundation models. These results indicate that ConceptCLIP effectively comprehends and responds to complex medical queries, demonstrating its potential as a decision-support tool in clinical environments where accurate medical decision-making is essential.
@@
ConceptCLIP improves the quality of medical report generation. Medical report generation is a critical task for summarizing complex medical information. As illustrated in Fig. 3 (a), the pre-trained image encoder extracts image features, which are concatenated with text embeddings and processed by a large language model to generate the final report (the Method section). Fig. 3 (d) and Extended Data Table 11 present the performance of ConceptCLIP on the medical report generation task using the MIMIC-CXR [17] and IU X-Ray [18] datasets. ConceptCLIP consistently outperforms or achieves competitive performance with other models on most metrics for these datasets. Notably, in MIMIC-CXR, ConceptCLIP achieves a Macro F1 score of 26.38% (95% CI: 25.37% - 27.41%), surpassing the second-best result by 4.28% ($P<0.001$), indicating its superior ability to generate accurate and coherent medical reports. This performance suggests that ConceptCLIP effectively understands and summarizes complex medical information, providing valuable support in clinical documentation. However, in IU-X-Ray, the Micro Precision, Recall, and F1 scores of ConceptCLIP were not as high as BiomedCLIP. This discrepancy is attributed to the imbalance in the IU X-Ray dataset, where most cases are normal. A model that consistently outputs normal results can achieve high Micro Precision, Recall, and F1 scores, indicating a bias in the dataset that does not fully reflect a model’s performance.
@@
ConceptCLIP enables precise pathology whole-slide image analysis. Whole-slide image analysis is a vital task for enhancing diagnostic accuracy in histopathology. As depicted in Fig. 3 (a), a whole-slide image is divided into patches, each of which is processed by the pre-trained vision encoder. The extracted visual features are aggregated for downstream classification tasks. As shown in Fig. 3 (e) and Extended Data Table 12, ConceptCLIP demonstrates exceptional performance across various tasks, including cancer diagnosis, mutation prediction, and survival prediction. For instance, in the cancer diagnosis task on the BRACS-3 dataset, ConceptCLIP achieves an AUC score of 91.65% (95% CI: 87.18% - 96.12%), outperforming all other models ($P<0.001$). Similarly, in the survival prediction task for LUSC, ConceptCLIP records the highest C-index of 59.66% (95% CI: 50.31% - 69.01%). These results highlight ConceptCLIP’s capability to process and analyze pathology images of a humongous size, making it a promising tool for enhancing diagnostic accuracy in histopathology.